A monatomic ideal gas (γ =1.67) is compressed adiabatically to one-twentieth its initial volume with P1 = 1.00 atm. What is the final pressure? Assuming T1=300K, What is T2(final temperature)?
Answer #1

given,

V2 = V1/ 20

now az per adiabatic equ

P1 V1^ Y = P2 V2^Y

P2 = P1 ( V1/V2) ^Y

P2 = 1 * (20)^1.67

P2 = 148.84 atm

=======

now,

P1 V1 / T1 = P2 V2 / T2

1 * 20 / 300 = 148.8 / T2

T2 = 2232.6 k

=======

Comment before rate in case any doubt, will reply for sure.. goodluck